# Global weak solutions for the inverse mean curvature flow in the Heisenberg group

### v2

@@ Abstract.
We consider the inverse mean curvature flow (IMCF) in the Heisenberg group $(\mathbb{H}^{n},d_{\varepsilon})$, where $d_{\varepsilon}$ is distance associated to either $|\cdot|_{\varepsilon}$, $\varepsilon>0$, the natural family of left-invariant Riemannian metrics, or with their sub-Riemannian counterparts for $\varepsilon=0$. For $\Omega\subseteq\mathbb{H}^{n}$ an open set with smooth boundary $\Sigma_{0}=\partial\Omega$ satisfying a uniform exterior gauge-ball condition and bounded complement we show existence of a global weak IMCF of generalized hypersurfaces $\{\Sigma^{\varepsilon}_{s}\}_{s\geq 0}\subseteq\mathbb{H}^{n}$ which is proper. Here, both in the Riemannian and in the sub-Riemannian setting, we adopt the weak formulation introduced by Huisken and Ilmanen in [28], following the approach in [34] due to Moser and based on the the link between IMCF and $p$-harmonic functions.
@@
E.V. is a member of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). A.P. is partially funded by PRIN project 2022PJ9EFL “Geometric Measure Theory: Structure of Singular Measures, Regularity Theory and Applications in the Calculus of Variations”. E.V. is partially funded by PRIN project 2022R537CS “$NO^{3}$ Nodal Optimization, NOnlinear elliptic equations, NOnlocal geometric problems, with a focus on regularity” and by the Indam-GNAMPA project “Esistenza, unicità e regolarità di soluzioni per problemi singolari”.
@@ 1. Introduction
(1.1) $\left\{\begin{array}[]{rl}{\rm div}_{0}\left(\dfrac{\nabla_{0}u}{|\nabla_{0}u|_{0}}\right)=|\nabla_{0}u|_{0}&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{array}\right.$
[…]
@@
where $\mathcal{H}_{H}\neq 0$ at a point $(p,s)\in\Sigma\times(0,S)$ is the horizontal mean curvature vector of the hypersurface $\Sigma_{s}=F(\Sigma,s)$ at the point $F(p,s)$. In case of regular hypersurfaces, if a smooth function $u:\overline{\Omega}\to[0,\infty)$ with $\nabla_{0}u\neq 0$ in $\overline{\Omega}$ satisfies (1.1) then the family $\{\Sigma^{\varepsilon}_{s}\}_{s\in[0,S)}$, with $\Sigma_{s}:=\{u=s\}$, $s\in[0,S)$ and $S>0$ small enough, is a smooth horizontal IMCF, with $\Sigma_{s}=F(\Sigma,s)$ and the map $F$ being the flow map associated to $\frac{\nabla_{0}u}{|\nabla_{0}u|_{0}^{2}}$. In particular, when the initial surface is $\Sigma_{0}=\partial B_{R_{0}}(g_{0})$ then according to (1.2) we have $\Sigma_{s}=\partial B_{R_{s}}(g_{0})$, where ${R_{s}=R_{0}e^{\frac{s}{Q-1}}}$ and $s\in[0,\infty)$, and indeed (1.1) is satisfied at least wherever $\nabla_{0}U\neq 0$, i.e., away from the center of the group.
@@
(1.3) $\left\{\begin{array}[]{rl}{\rm div}_{\varepsilon}\left(\dfrac{\nabla_{\varepsilon}u}{|\nabla_{\varepsilon}u|_{\varepsilon}}\right)=|\nabla_{\varepsilon}u|_{\varepsilon}&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{array}\right.$
where $\rm{div}_{\varepsilon}$ and $\nabla_{\varepsilon}u$ are computed w.r.to the standard left invariant Riemannian metrics $|\cdot|_{\varepsilon}$ and the equation (1.3) give the analogous level set formulation for the corresponding IMCFs.
@@
among functions $w$ which are compactly supported perturbations of $u$, i.e., $\overline{\{u\neq w\}}\subset K$ and $K\subset\Omega$ any compact set. Here and throughout the paper, all the volume integrals are computed w.r.to the Lebesgue measure $\mathcal{L}^{2n+1}$ in $\mathbb{R}^{2n+1}\simeq\mathbb{H}^{n}$ (i.e., up to a factor the unique Haar measure on $\mathbb{H}^{n}$), also denoted sometimes by ||$\cdot$|| ⋅ | in the sequel, and it will be deliberately omitted in all the integrals.
@@
In order to deal with (1.1)-(1.3) we adapt the very clever approach introduced in the important paper [34] for the Euclidean case, see also [35], and then fruitfully exploited in the Riemannian and Finsler case, see e.g. [22, 23, 32, 40, 16, 6]. Thus, for fixed $\varepsilon\in(0,1]$ we are are going to construct solutions $u^{\varepsilon}:\overline{\Omega}\subseteq\mathbb{H}^{n}\to{\mathbb{R}}$ of (1.3) as limits as $p\downarrow 1$ of solutions $u^{\varepsilon}_{p}:\overline{\Omega}\subseteq\mathbb{H}^{n}\to{\mathbb{R}}$ to the regularized problems
(1.6a) ${\rm div}_{\varepsilon}\left(|\nabla_{\varepsilon}u^{\varepsilon}_{p}|_{\varepsilon}^{p-2}\nabla_{\varepsilon}{u^{\varepsilon}_{p}}\right)=|\nabla_{\varepsilon}u^{\varepsilon}_{p}|_{\varepsilon}^{p}$ in $\quad\textrm{in }\Omega,$
(1.6b) $u^{\varepsilon}_{p}=0$ on $\quad\textrm{on }\partial\Omega.$
@@
As it will be shown below, for $0\leq\varepsilon\leq 1$ and $1<p<Q$ problems (1.6a) admit solutions $u_{p}^{\varepsilon}$ in $C(\overline{\Omega})\cap C^{1,\beta}(\Omega)$ where the validity of the equations is a consequence of (but actually equivalent to, see Lemma 7.3) the local minimality property analogous to (1.5) for the energy functionals
@@
As observed in [34], for $p>1$ and $0\leq\varepsilon\leq 1$ the substitution
@@
(1.9a) ${\rm div}_{\varepsilon}\left(|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{p-2}\nabla_{\varepsilon}v^{\varepsilon}_{p}\right)=0$ in $\quad\textrm{in }\Omega,$
(1.9b) $v^{\varepsilon}_{p}=1$ on $\quad\textrm{on }\partial\Omega.$
@@
Concerning problem (1.9a)-(1.9b) we will only consider for $1<p<Q$ finite energy solutions, i.e., those minimizing the energy functionals
@@
Here and in the sequel $\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$ denotes for any $\varepsilon\in[0,1]$ the homogeneous Sobolev space of functions in $u\in L^{p^{*}}(\mathbb{H}^{n})$, $p^{*}=\frac{pQ}{Q-p}$, with weak gradient $\nabla_{\varepsilon}u$ such that $|\nabla_{\varepsilon}u|_{\varepsilon}\in L^{p}(\mathbb{H}^{n})$, endowed with the norm $\|u\|_{\dot{W}_{\varepsilon}^{1,p}}=\|\,|\nabla_{\varepsilon}u|_{\varepsilon}\|_{L^{p}}$, the case $\varepsilon=0$ corresponding to the horizontal Sobolev space $H\dot{W}^{1,p}(\mathbb{H}^{n})$ first introduced by Folland and Stein. Clearly such spaces are well defined because of the horizontal Sobolev inequality corresponding to the embedding $H\dot{W}^{1,p}(\mathbb{H}^{n})\subset L^{p^{*}}(\mathbb{H}^{n})$ and actually valid for $1\leq p<Q$ (see (2.4) below).
Note that such optimal functions solving (1.9a)-(1.9b) are nothing but the $p-$capacitary potentials of $\Omega^{c}$. For $1<p<Q$ and $\varepsilon\geq 0$ existence and uniqueness for this minimization problem is a standard consequence of the direct method in the Calculus of Variations for the coercive weakly l.s.c. functional (1.10) on the weakly closed convex subset (1.11) of the reflexive space $\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$. For $\varepsilon>0$ the $C^{1,\beta}$-regularity up to the boundary is nowadays classical in the regularity theory for $p$-harmonic functions on Riemannian manifolds (see, e.g., [17] and [31]), with exponent $\beta=\beta(p,\varepsilon)\in(0,1)$. Interior regularity is also known in the sub-Riemannian case $\varepsilon=0$, see, e.g., [37], so that $p-$capacitary potentials actually belong to the space $C(\overline{\Omega})$, where continuity up to the boundary is a consequence of the exterior gauge ball condition ($\mathbf{HP_{\Omega}}$), and with horizontal gradient in $C^{0,\beta^{\prime}}(\Omega)$ i for some $\beta^{\prime}\in(0,1)$. On the other hand, continuity up to the boundary of the horizontal gradient is at present not available for (1.9a) when $\varepsilon=0$ and this is indeed the major technical reason for us to address existence theory for (1.1) through the analysis of the two-parameter family of problems (1.6a)-(1.6b) together with their $p-$harmonic counterparts (1.9a)-(1.9b).
@@
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying the exterior uniform gauge-ball condition ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For any $1<p<Q$ and $\varepsilon\in[0,1]$ let $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the unique finite energy solution to (1.9a) corresponding to the minimizer of (1.10) in the set (1.11). Then $v^{\varepsilon}_{p}\in C^{1}(\overline{\Omega})$ for any $\varepsilon\in(0,1]$, it is strictly positive in $\overline{\Omega}$ and there exists a constant $C>0$ depending only on $R_{0}$ and $Q$ such that in $\overline{\Omega}$ we have
@@
Moreover, for $1<p<Q$ and $\varepsilon=0$ the corresponding finite energy solution $v^{0}_{p}\in\dot{W}^{1,p}_{1,0}(\Omega)$ to (1.9a) satisfies $v^{0}_{p}\in C(\overline{\Omega})$, $\nabla_{0}v^{0}_{p}\in C(\Omega;\mathbb{R}^{2n})$ and (1.12) holds in $\Omega$ with the same $C>0$ above.
In view of (1.8), as a straightforward consequence of Theorem 1.2 one gets the following uniform gradient bound for solutions to (1.6a)-(1.6b).
@@
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement such that $\Omega$ satisfying assuption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For $1<p<Q$ and $\varepsilon\in[0,1]$ let $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the solution to (1.9a) corresponding to the unique minimizers of (1.10) in the set (1.11) and $u_{p}^{\varepsilon}$ given by (1.8). Then there exists $C>0$ depending only on $R_{0}$ and $Q$ such that
@@
As already suggested above, the case $\varepsilon=0$ in the theorem is obtained by a limiting argument from the case $\varepsilon\in(0,1]$ for which boundary regularity is available, as the minimizers $\{v^{\varepsilon}_{p}\}$ are strongly convergent in $\dot{HW}^{1,p}(\mathbb{H}^{n})$ to $v^{0}_{p}$ as $\varepsilon\to 0$. Concerning the latter, here the strategy in the proof of Theorem 1.2 and in turn Corollary 1.3 is similar to the one adopted in [34], i.e., one constructs barriers to obtain the bound (1.12) for $v^{\varepsilon}_{p}$ at the boundary $\partial\Omega$ and then propagate this bound in the interior by a suitable use of the maximum principle and with some extra care since $\Omega$ is not bounded (see the remarks at the beginning of Section 6 for a more detailed discussion on this aspect).
Here, in constrast with [34] (as well as [30] and [15]), instead of working with $u^{\varepsilon}_{p}$ solving (1.6a) and applying the maximum principle on the equation satisfyied by $|\nabla_{\varepsilon}u^{\varepsilon}_{p}|^{2}$ we find more convenient to work directly with $v^{\varepsilon}_{p}$ and the corresponding quantity $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{2}-C^{2}(v^{\varepsilon}_{p})^{2}$ with the value of $C$ already obtained for the gradient estimate at the boundary. In addition, compared to [34] both the steps mentioned above turn out to be more difficult in the present case. Indeed, explicit solutions in terms of negative powers of the Korányi norm are available only for $\varepsilon=0$, whence, still under assumption ($\mathbf{HP_{\Omega}}$), a suitable perturbation of them must be used as barriers along $\partial\Omega$ in the full range $\varepsilon\in[0,1]$. Furthermore, the maximum principle argument alluded above actually does not seem to apply directly to $|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{2}-C^{2}(v^{\varepsilon}_{p})^{2}$, so that we have to estimate separately first the vertical derivative $|T_{\varepsilon}v_{p}^{\varepsilon}|^{2}$ and then $|\nabla_{0}v^{\varepsilon}_{p}|_{0}^{2}$, by exploiting the explicit structure in $\mathbb{H}^{n}$ of the Ricci tensor $Ric_{\varepsilon}$ in connection with a key computation adapted from [30] involving the linearized operator (6.8) of (1.9a) at $v^{\varepsilon}_{p}$.
@@
As for those in [34], both the steps our proof rely in an essential way on the $C^{1}$-boundary regularity of $p$-capacitary potentials on Riemannian manifolds, as they need to be differentiable at the boundary for the first step and must be of controlled size in tubular neighborhood of $\partial\Omega$ in order to obtain the second step (by running a contradiction argument well in the interior where even higher regularity is available and used). To our knowledge such boundary regularity property is not yet available in the sub-Riemannian case, therefore we have no direct proof of the theorem in the case $\varepsilon=0$ but we can only infer the same bound from (1.12) by letting $\varepsilon\to 0$, as already announced above. The same lack of boundary regularity affects the argument used for the recent existence result [15, Proposition 5.3] for the boundary value problem (1.1) in $\mathbb{H}^{1}$, so that their proof seems to contain a serious gap and our aim here is to provide an alternative approach to such existence property for IMCF on $\mathbb{H}^{n}$ through the Riemannian approximation.
@@
In the Riemannian case, local gradient bounds for $p$-harmonic functions similar to (1.13) were first obtained in the literature in [30] on complete manifolds under lower bounds on the sectional curvature (see also [39] and [32] for similar results under the weaker assumption of lower bounds on the Ricci curvature). However, notice that along the vertical vector field $T_{\varepsilon}=\varepsilon\partial_{t}$ one has $|T_{\varepsilon}|_{\varepsilon}\equiv 1$ but $Ric_{\varepsilon}(T_{\varepsilon})\to-\infty$ as $\varepsilon\to 0$. Moreover, being uniform w.r.to $\varepsilon$, our global bounds here are clearly stronger and do not follow from previous results in the existing literature. On the other hand, it would be interesting to obtain directly purely sub-Riemannian local gradient estimates as done e.g. [3] in the linear case $p=2$.

### v3

@@ Abstract.
We consider the inverse mean curvature flow (IMCF) in the Heisenberg group $(\mathbb{H}^{n},d_{\varepsilon})$, where $d_{\varepsilon}$ is distance associated to either $|\cdot|_{\varepsilon}$, $\varepsilon>0$, the natural family of left-invariant Riemannian metrics, or with their sub-Riemannian counterparts for $\varepsilon=0$. For $\Omega\subseteq\mathbb{H}^{n}$ an open set with smooth boundary $\Sigma_{0}=\partial\Omega$ satisfying a uniform exterior gauge-ball condition and bounded complement we show existence of a global weak IMCF of generalized hypersurfaces $\{\Sigma^{\varepsilon}_{s}\}_{s\geq 0}\subseteq\mathbb{H}^{n}$ which are level sets of a proper globally Lipschitz function with logarithmic growth at infinity. Here, both in the Riemannian and in the sub-Riemannian setting, we adopt the weak formulation introduced by Huisken and Ilmanen in [29], following the approach in [36] due to Moser and based on the the link between IMCF and $p$-harmonic functions.
@@ 1. Introduction
(1.1) $\left\{\begin{array}[]{ll}{\rm div}_{0}\left(\dfrac{\nabla_{0}u}{|\nabla_{0}u|_{0}}\right)=|\nabla_{0}u|_{0}&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{array}\right$.
where $\rm{div}_{0}$ is the horizontal divergence, $\nabla_{0}u$ is the horizontal gradient of $u$ and its length is computed w.r.to the standard sub-Riemannian metric $|\cdot|_{0}$ on $\mathbb{H}^{n}$. In the model case of the complement of a gauge-ball, i.e., $\Omega=\mathbb{H}^{n}\setminus\overline{B_{R_{0}}(g_{0})}$ for some $g_{0}\in\mathbb{H}^{n}$ and $R_{0}>0$, an explicit solution of (1.1) can be written in terms of the Korányi norm $\|\cdot\|$ as
@@
where $\mathcal{H}_{H}\neq 0$ at a point $(p,s)\in\Sigma\times(0,S)$ is the horizontal mean curvature vector of the hypersurface $\Sigma_{s}=F(\Sigma,s)$ at the point $F(p,s)$. In case of regular hypersurfaces, if a smooth function $u:\overline{\Omega}\to[0,\infty)$ with $\nabla_{0}u\neq 0$ in $\overline{\Omega}$ satisfies (1.1) then the family $\{\Sigma_{s}\}_{s\in[0,S)}$, with $\Sigma_{s}:=\{u=s\}$, $s\in[0,S)$ and $S>0$ small enough, is a smooth horizontal IMCF, with $\Sigma_{s}=F(\Sigma,s)$ and the map $F$ being the flow map associated to the vector field $\frac{\nabla_{0}u}{|\nabla_{0}u|_{0}^{2}}$. In particular, when the initial surface is $\Sigma_{0}=\partial B_{R_{0}}(g_{0})$ then according to (1.2) we have $\Sigma_{s}=\partial B_{R(s)}(g_{0})$, where ${R(s)=R_{0}e^{\frac{s}{Q-1}}}$ and $s\in[0,\infty)$, and indeed (1.1) is satisfied at least wherever $\nabla_{0}U\neq 0$, i.e., at those points of the corresponding level sets which are not characteristic.
@@
(1.3) $\left\{\begin{array}[]{ll}{\rm div}_{\varepsilon}\left(\dfrac{\nabla_{\varepsilon}u}{|\nabla_{\varepsilon}u|_{\varepsilon}}\right)=|\nabla_{\varepsilon}u|_{\varepsilon}&\textrm{in }\Omega,\\ u=0&\textrm{on }\partial\Omega,\end{array}\right$.
where $\rm{div}_{\varepsilon}$ and $\nabla_{\varepsilon}u$ are computed w.r.to the standard left invariant Riemannian metrics $|\cdot|_{\varepsilon}$ and the equation (1.3) give the analogous level set formulation for the corresponding IMCFs. Here the actual interest in (1.3) is twofold: on the one hand viewing it as an $\varepsilon$-regularized version of (1.1) but on the other hand even in the case $\varepsilon=1$ a solution to (1.3) giving a family of hypersurfaces expanding by IMCF in a Riemannian manifold with an intrinsically sub-Riemannian nature at large scales (indeed, as it is well-known, the sub-Riemannian space being the tangent cone at infinity of the Riemannian one under Gromov-Hausdorff convergence of its rescaled copies under group dilations), a feature which deeply affects the large time behaviour of the flow.
@@
among functions $w$ which are compactly supported perturbations of $u$, i.e., $\overline{\{u\neq w\}}\subset K$ and $K\subset\Omega$ any compact set. Here and throughout the paper, all the volume integrals are computed w.r.to the Lebesgue measure $\mathcal{L}^{2n+1}$ in $\mathbb{R}^{2n+1}\simeq\mathbb{H}^{n}$ (i.e., up to a factor the unique Haar measure on $\mathbb{H}^{n}$), also denoted sometimes by $|\cdot|$ in the sequel, and it will be deliberately omitted in all the integrals.
@@
In order to deal with (1.1)-(1.3) we adapt the very clever approach introduced in the important paper [36] for the Euclidean case (see also [37]), and then fruitfully exploited both in the Riemannian and in Finsler geometries, see e.g. [23, 24, 33, 43, 17, 6]. Thus, for fixed $\varepsilon\in(0,1]$ we are are going to construct solutions $u^{\varepsilon}:\overline{\Omega}\subseteq\mathbb{H}^{n}\to{\mathbb{R}}$ of (1.3) as limits as $p\downarrow 1$ of solutions $u^{\varepsilon}_{p}:\overline{\Omega}\subseteq\mathbb{H}^{n}\to{\mathbb{R}}$ to the regularized problems
(1.6a) ${\rm div}_{\varepsilon}\left(|\nabla_{\varepsilon}u^{\varepsilon}_{p}|_{\varepsilon}^{p-2}\nabla_{\varepsilon}{u^{\varepsilon}_{p}}\right)=|\nabla_{\varepsilon}u^{\varepsilon}_{p}|_{\varepsilon}^{p}$ $\quad\textrm{in }\Omega,$
(1.6b) $u^{\varepsilon}_{p}=0$ $\quad\textrm{on }\partial\Omega.$
@@
As it will be shown below, for $0<\varepsilon\leq 1$ and $1<p<Q$ problems (1.6a) admit solutions $u_{p}^{\varepsilon}$ in $C(\overline{\Omega})\cap C^{1,\beta}(\Omega)$ where the validity of the equations is a consequence of (but actually equivalent to, see Lemma 8.4) the local minimality property analogous to (1.5) for the energy functionals
@@
According to the beautiful observation made in [36], for $p>1$ and $0\leq\varepsilon\leq 1$ the substitution
@@
(1.9a) ${\rm div}_{\varepsilon}\left(|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{p-2}\nabla_{\varepsilon}v^{\varepsilon}_{p}\right)=0$ $\quad\textrm{in }\Omega,$
(1.9b) $v^{\varepsilon}_{p}=1$ $\quad\textrm{on }\partial\Omega.$
@@
Concerning problem (1.9a)-(1.9b), in the present paper we will only consider for $1<p<Q$ finite energy solutions, i.e., those minimizing the energy functionals
@@
Here and in the sequel $\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$ denotes for any $\varepsilon\in[0,1]$ the homogeneous Sobolev space of functions in $u\in L^{p^{*}}(\mathbb{H}^{n})$, $p^{*}=\frac{pQ}{Q-p}$, with distributional gradient $\nabla_{\varepsilon}u$ such that $|\nabla_{\varepsilon}u|_{\varepsilon}\in L^{p}(\mathbb{H}^{n})$, endowed with the norm $\|u\|_{\dot{W}_{\varepsilon}^{1,p}}=\|\,|\nabla_{\varepsilon}u|_{\varepsilon}\|_{L^{p}}$, the case $\varepsilon=0$ corresponding to the homogeneous horizontal Sobolev space $H\dot{W}^{1,p}(\mathbb{H}^{n})$ first introduced by Folland and Stein. Clearly such spaces are well defined because of the horizontal Sobolev inequality corresponding to the embedding $H\dot{W}^{1,p}(\mathbb{H}^{n})\subset L^{p^{*}}(\mathbb{H}^{n})$ and actually valid for $1\leq p<Q$ (see (2.6) below).
Note that such optimal functions solving (1.9a)-(1.9b) are nothing but the $p$-capacitary potentials of $\Omega^{c}$. For $1<p<Q$ and $\varepsilon\geq 0$ existence and uniqueness for this minimization problem is a standard consequence of the direct method in the Calculus of Variations for the coercive weakly l.s.c. functional (1.10) on the weakly closed convex subset (1.11) of the reflexive space $\dot{W}_{\varepsilon}^{1,p}(\mathbb{H}^{n})$. For $\varepsilon>0$ the $C^{1,\beta}$-regularity up to the boundary is nowadays classical in the regularity theory for $p$-harmonic functions on Riemannian manifolds (see, e.g., [18] and [32]), with exponent $\beta=\beta(p,\varepsilon)\in(0,1)$. Interior regularity is also known in the sub-Riemannian case $\varepsilon=0$, see, e.g., [39], so that $p$-capacitary potentials actually belong to the space $C(\overline{\Omega})$, where continuity up to the boundary is a consequence of the exterior gauge ball condition ($\mathbf{HP_{\Omega}}$), and with horizontal gradient in $C^{0,\beta^{\prime}}(\Omega)$ for some $\beta^{\prime}\in(0,1)$. On the other hand, continuity up to the boundary of the horizontal gradient is at present not available for (1.9a) when $\varepsilon=0$ and this is indeed the major technical reason for us to address existence theory for (1.1) through the analysis of the two-parameter family of problems (1.6a)-(1.6b) together with their $p$-harmonic counterparts (1.9a)-(1.9b).
@@
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying the exterior uniform gauge-ball condition ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For any $1<p\leq 2$ and $\varepsilon\in[0,1]$ let $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the unique finite energy solution to (1.9a) corresponding to the minimizer of (1.10) in the set (1.11). Then $v^{\varepsilon}_{p}\in C^{1}(\overline{\Omega})$ for any $\varepsilon\in(0,1]$, it is strictly positive in $\overline{\Omega}$ and there exists a constant $C>0$ depending only on $R_{0}$ and $Q$ such that in $\overline{\Omega}$ we have
@@
Moreover, for $1<p\leq 2$ and $\varepsilon=0$ the corresponding finite energy solution $v^{0}_{p}\in\dot{W}^{1,p}_{1,0}(\Omega)$ to (1.9a) satisfies $v^{0}_{p}\in C(\overline{\Omega})$, $\nabla_{0}v^{0}_{p}\in C(\Omega;\mathbb{R}^{2n})$ and (1.12) holds in $\Omega$ with the same $C>0$ above.
In view of (1.8), as a straightforward consequence of Theorem 1.2 (together with Proposition 8.2) one gets the following uniform gradient bound for solutions to (1.6a)-(1.6b).
@@
Let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement such that $\Omega$ satisfying assumption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$. For $1<p\leq 2$ and $\varepsilon\in[0,1]$ let $v^{\varepsilon}_{p}\in\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ the solution to (1.9a) corresponding to the unique minimizers of (1.10) in the set (1.11) and $u_{p}^{\varepsilon}$ given in turn by (1.8). Then there exists $C>0$ depending only on $R_{0}$ and $Q$ such that
@@
As a consequence, $u_{p}^{\varepsilon}\in Lip_{\varepsilon}(\mathbb{H}^{n})$ with $d_{\varepsilon}$-Lipschitz constant satisfying the uniform bound $Lip_{\varepsilon}(u_{p}^{\varepsilon})\leq C$.
@@
As already suggested above, the case $\varepsilon=0$ in the theorem is obtained by a limiting argument from the case $\varepsilon\in(0,1]$ for which boundary regularity is available, as the minimizers $\{v^{\varepsilon}_{p}\}$ are strongly convergent in $\dot{HW}^{1,p}(\mathbb{H}^{n})$ to $v^{0}_{p}$ as $\varepsilon\to 0$. Concerning the latter, here the strategy in the proof of Theorem 1.2 and in turn Corollary 1.3 is similar in the spirit to the one adopted in [36], i.e., one constructs barriers to obtain the bound (1.12) for $v^{\varepsilon}_{p}$ at the boundary $\partial\Omega$ and then propagate this bound in the interior by a suitable use of the maximum principle (with some extra care since $\Omega$ is not bounded). However, compared to [36] both the steps mentioned above turn out to be more difficult in the present case. Indeed, explicit solutions in terms of negative powers of the Korányi norm are available only for $\varepsilon=0$, whence, still under assumption ($\mathbf{HP_{\Omega}}$), a suitable perturbation of them must be used as barriers along $\partial\Omega$ in the full range $\varepsilon\in[0,1]$. On the other hand, as we explain below, propagating the gradient estimates to the interior turns out to be highly nontrivial in the present case.
Actually, in contrast with [36] (as well as [31], [16] and [33]), we are not able to derive uniform interior gradient estimates working directly with $u^{\varepsilon}_{p}$ solving (1.6a) and applying the maximum principle on the equation satisfied by $|\nabla_{\varepsilon}u^{\varepsilon}_{p}|^{2}$. Indeed, in contrast with the aforementioned papers where some lower bounds and asymptotic nonnegativity of the curvature at infinity is assumed, on $\mathbb{H}^{n}$ along the horizontal distribution one has $Ric_{\varepsilon}(\cdot)\equiv-\frac{1}{2\varepsilon^{2}}|\cdot|_{\varepsilon}^{2}$ everywhere. Moreover, $Ric_{\varepsilon}(\cdot)\to$-$\infty$ as $\varepsilon\to 0$ in any horizontal direction, whereas $Ric_{\varepsilon}(\cdot)\to+\infty$ along the vertical direction, which forces us to devise a completely new approach adapted to the intrinsic structure of $\mathbb{H}^{n}$ to infer the global interior gradient estimates.
Here we find more convenient to work directly with $v^{\varepsilon}_{p}$ as a solution to (1.9a) and to rely on the Bochner method. Thus we first consider the corresponding quantity $(p-1)^{2}(Tv^{\varepsilon}_{p})^{2}-C^{2}(v^{\varepsilon}_{p})^{2}$, with the value $C=C(R_{0})>0$ ensuring nonpositivity of such a quantity at the boundary by a barrier argument, and then we infer non positivity also in the interior by a simple maximum principle argument using the elliptic operator $L_{v^{\varepsilon}_{p}}$ in (6.1), i.e., the linearized $p$-Laplace operator obtained from (1.9a) at $v^{\varepsilon}_{p}$. Once this Cheng-Yau’s type inequality for the vertical derivative $T=\partial_{t}$ is obtained, we next obtain the analugue for the full gradient, i.e., (1.12), by a similar but more involved maximum principle argument. Here the heart of the matter is to establish an improved Kato-type inequality for the squared norm of the rough Hessian $\left(\overline{D}^{2}v^{\varepsilon}_{p}\right)_{i,j}=E_{j}(E_{i}v^{\varepsilon}_{p})$ which is similar to those improved Kato inequality for the Frobenius norm of the covariant Hessian for $p$-harmonic functions or $p$-harmonic maps as already established in [40], [14] and [34]. In turn this improved inequality yields a Bochner inequality for $L_{v^{\varepsilon}_{p}}\left(|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{2}+(Tv^{\varepsilon}_{p})^{2}\right)$, the latter crucially relying on the explicit structure the Ricci tensor $Ric_{\varepsilon}$ in $\mathbb{H}^{n}$ in combination with a key computation adapted from [31] giving an explicit identity for $L_{v^{\varepsilon}_{p}}\left(|\nabla_{\varepsilon}v^{\varepsilon}_{p}|_{\varepsilon}^{2}\right)$ alone.
@@
As for those in [36], both the steps our proof rely in an essential way on the $C^{1}$-boundary regularity of $p$-capacitary potentials on Riemannian manifolds, as they need to be differentiable at the boundary for the first step and must be of controlled size in tubular neighborhood of $\partial\Omega$ in order to obtain the second step (by running a contradiction argument well in the interior where even higher regularity is available and used). To our knowledge such boundary regularity property is not yet available in the sub-Riemannian case, therefore we have no direct proof of the theorem in the case $\varepsilon=0$ but we can only infer the same bound from (1.12) by letting $\varepsilon\to 0$, as already announced above. The same lack of boundary regularity affects the argument used for the recent existence result [16, Proposition 5.3] for the boundary value problem (1.1) in $\mathbb{H}^{1}$, so that their proof seems to contain a serious gap and our aim here is to provide an alternative approach to such existence property for IMCF on $\mathbb{H}^{n}$ through the Riemannian approximation.
@@
Notice that in the Riemannian case local gradient bounds for $p$-harmonic functions similar to (1.13) were first obtained in the literature in [31] on complete manifolds under lower bounds on the sectional curvature (see also [42] and [33] for similar results under the weaker assumption of lower bounds on the Ricci curvature). However, as we have already recalled that along the horizontal distribution of $\mathbb{H}^{n}$ one has $Ric_{\varepsilon}(\cdot)\equiv-\frac{1}{2\varepsilon^{2}}|\cdot|_{\varepsilon}^{2}\to$-$\infty$ as $\varepsilon\to 0$ it seems difficult to obtain a local Cheng-Yau’s inequality in the sub-Riemannian case through the Riemannian approximation. Thus, although being global in nature but being uniform w.r.to $\varepsilon$, our gradient bounds here are clearly stronger and they do not follow from previous results in the existing literature. However, as an alternative route to the limiting argument used here, it would be interesting to obtain directly a sub-Riemannian local gradient estimates as done e.g. [3] in the case of harmonic functions on Carnot groups, although the case of non quadratic energies seems at present awkward.